# FocusFlow: 3D Gaze-Depth Interaction in Virtual Reality Leveraging Active Visual Depth Manipulation

### v2

@@ Abstract.
Gaze interaction presents a promising avenue in Virtual Reality (VR) due to its intuitive and efficient user experience. Yet, the depth control inherent in our visual system remains underutilized in current methods. In this study, we introduce FocusFlow, a hands-free interaction method that capitalizes on human visual depth perception within the 3D scenes of Virtual Reality. We first develop a binocular visual depth detection algorithm to understand eye input characteristics. We then propose a layer-based user interface and introduce the concept of “Virtual Window” that offers an intuitive and robust gaze-depth VR interaction, despite the constraints of visual depth accuracy and precision spatially at further distances. Finally, to help novice users actively manipulate their visual depth, we propose two learning strategies that use different visual cues to help users master visual depth control. Our user studies on 24 participants demonstrate the usability of our proposed virtual window concept as a gaze-depth interaction method. In addition, our findings reveal that the user experience can be enhanced through an effective learning process with adaptive visual cues, helping users to develop muscle memory for this brand-new input mechanism. We conclude the paper by discussing strategies to optimize learning and potential research topics of gaze-depth interaction.
gaze interaction, visual depth, virtual reality, user interface design
journalyear: 2024copyright: acmlicensedconference: Proceedings of the CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, HI, USAbooktitle: Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613904.3642589isbn: 979-8-4007-0330-0/24/05ccs: Human-centered computing HCI design and evaluation methodsccs: Human-centered computing Virtual realityccs: Human-centered computing Interaction design theory, concepts and paradigms
@@ 1. Introduction
Gaze interaction has gained popularity as an input method for 3D interaction in mixed reality (XR) headsets (Sidenmark et al., 2021; Lu et al., 2021a; Adhanom et al., 2023), where the eye movements and gestures would be used as an input. Current gaze-based XR interaction methods often rely on gaze as a means of pointing, supplemented by other inputs such as hand gestures (Yu et al., 2021), head movement (Sidenmark et al., 2021), or dwell time (Penkar et al., 2012; Piumsomboon et al., 2017) for selection confirmation to avoid the so-called ”Midas Touch” problem (Velichkovsky et al., 1997) – the problem of distinguishing intentional gaze inputs from involuntary fixations as the users perceive the scene. However, these works primarily utilize the direction of gaze and overlook the valuable gaze depth information, representing an additional free input dimension along the z-axis (Figure 1). To see how, imagine a window through which you can see scenery outside. By bringing your gaze focus closer to the window, you blur out the distant scenery and are able to perceive dust particles on the window surface instead. Another example is autostereograms (Tyler and Clarke, 1990), where slightly adjusting your visual depth allows you to perceive a 3D image (see example in Appendix A). In both of these examples, users actively manipulate their visual depth to select which object to observe. In this research, we aim to leverage this natural eye behavior and convert it into an intuitive and learnable interaction input in Virtual Reality (VR).
@@
More recent works on gaze-based VR/AR interaction demonstrate the great potential of visual depth as an interaction input to solve the Midas touch problem. These methods either guide the user to look at physical or virtual objects at different depths (Ahn et al., 2020; Pai et al., 2016; Wang et al., 2022a, b; Sidenmark et al., 2023) or rely on voluntary eye convergence and divergence (Kirst and Bulling, 2016; Kudo et al., 2013) by asking the users focusing on the nose or imagining to fixate on some point behind the display plane. However, these works lack an intuitive and systematic User Interface (UI) design to guide the users in manipulating their visual depth, leading to limited application scenarios and potentially user frustration. In addition, most of these works lack a thorough user study to evaluate usability and efficiency of this interaction modality.
@@
In particular, leveraging visual depth as an interaction input is challenging due to several practical factors. First, the natural eye movement is inherently unstable, resulting in noisy visual depth estimates. Moreover, the eye tracking modules in the headsets suffer from limited frame rate and jitters, adding extra noise to depth estimates. So, a usable interaction mechanism is required that is robots against imprecise eye input signal. Second, user’s depth fixation naturally relies on the presence of objects. So, it may feel counter-intuitive if the user is asked to transition to a depth without any visual referent. So, it is crucial to ensure that novice users can intuitively and effortlessly learn and use this interaction and a carefully designed learning procedure is needed to guide the users on how to master this new interaction mechanism. To tackle these challenges, we employed an empirical research approach. Initially, we devised a visual depth detection method for off-the-shelf VR headsets and conducted experiments to investigate the properties of visual depth input. Building upon the insights gained from these experiments, we subsequently developed our proposed user interface and incorporated a learning procedure with in-stages visual cues to help users adapt to the depth-control process.
As such, we present FocusFlow, a novel gaze-depth interaction approach that realizes the concept of “Virtual Window” (just like the examples mentioned above) by leveraging the 3D nature of immersive scenes and the human ability to actively shift between different visual depths. FocusFlow employs the eye trackers in VR headsets to estimate the user’s visual depth. It then defines a layer-based UI design with gaze interaction inputs. Layers are defined as transparent virtual windows at different depths that can be activated through voluntary change of visual depth to where the window is located. More specifically, FocusFlow uses the user’s gaze direction for selecting the object of interest for interaction and user’s visual depth to activate virtual windows that are otherwise invisible in the VR scene (shown in Figure 1). These virtual windows can offer different applications, from a quick preview to activating new control panels, or offering a zoom lens with more information about a selected object.
@@
To master this new interaction mechanism, we incorporate adaptive visual cues in the UI design and a learning process that allows users to establish a clear connection between the eye behavior and visual depth control, facilitating an intuitive and effortless interaction for novice users. The key intuition is that users can develop muscle memory for freely shifting between different focal depths even if empty of objects. We propose two learning strategies, in-stages learning, and adaptive learning. The in-stages learning uses a series of strong to weak visual cues at a certain depth (e.g. where the transparent virtual window is located) so the user can practice visual depth transitions by just moving her gaze from the target object to the cue. While these cues are effective in practicing gaze-depth control, they cause visual interference for the user during non-interaction times. The proposed adaptive learning process addresses this issue by adaptively changing the transparency of visual cues as the user gets closer to the intended interaction depth. The transparency level of the cue also serves as a feedback for the user to better control their focal depth changes. To help master this interaction, FocusFlow gradually reduces the guided depth range until reaching no visual cue.
@@
To assess the usability and learnability of our proposed interaction, we designed both quantitative and qualitative evaluations that take into account factors such as input efficiency, eye behavior, and cognitive load. The evaluation results from 24 participants prove that FocusFlow achieves high input efficiency and intuitive user experience on the hands-free selection tasks. Our design solves Midas Touch problem while retaining a natural interaction process. In addition, the result shows users can easily learn and master the gaze-depth method through our adaptive learning strategy. It is evident that the adaptive visual cues incorporated within our design substantially augmented users’ depth perception during the visual depth transition phase.
To the best of our knowledge, this work is the first to offer an intuitive and efficient gaze-depth interaction method with extensive evaluation. In summary, this presented work makes the following contributions:
@@
    We extensively analyze the characteristics of gaze-depth estimates and propose strategies for effectively utilizing this imprecise input information.
@@
    We develop a layer-based user interface for gaze-depth interaction by realizing the concept of a “virtual window”. This interface enables intuitive and robust hands-free VR interaction inputs while avoiding Midas Touch problem.
@@
    We presented a series of visual cues that provide users with depth perception to intuitively adjust their visual depth for interaction.
@@
    We present and compare two learning strategies, in-stages learning and adaptive learning, to guide novice users in adapting to depth control and gaze-depth interaction in virtual scenes. This learning process enables users with an effortless interaction by helping them develop muscle memory for this new interaction.
  5. (5)
    We conduct an user study and a case study to assess both the system’s performance and the user’s learning process. The results demonstrate the usability and learnability of our proposed system.
@@ 2. Related Work
In this section, we first summarize different eye behaviors and the corresponding input techniques which can be used for gaze interactions. Based on this, we then review the evolution of gaze-based interaction in VR and some representative methods. Finally, we discussed the existing gaze-depth interaction techniques to discover the research gap.
@@ 2.1. Eye Behaviors and Input Techniques
$2.$ Gaze-Based VR Interaction
@@
Gaze-depth-based Interaction. Gaze vergence is an emerging eye-input information that has been used for VR interaction. Compared to multi-modal gaze interaction methods that leverage handheld controllers or hand gestures, visual depth offers a completely hands-free interaction mechanism. Previous works have shown the feasibility of visual depth estimations in the VR headsets by using binocular disparity and stereoscopic vergence in VR headset (Hirzle et al., 2019; Vidal et al., 2014). The concept of visual depth as a new interaction input has explored in the previous works, either by defining a semi-transparent window at a different focal depth (Pai et al., 2016; Wang et al., 2022a, b), tracking voluntary vergence movements (Kirst and Bulling, 2016; Ahn et al., 2020), or matching the vergence changes with the depth changes of a moving objects (Sidenmark et al., 2023) in VR. However, all of these methods lack an end-to-end UI design to guide the users how to actively manipulate their visual depth. As a result, their users suffered from fatigue and difficulty to master this new interaction. In addition, they lack sufficient user studies to evaluate the usability and efficiency of this interaction mechanism in the context of different applications. This paper targets these limitations and proposes a layer-based UI design to leverage active visual depth manipulation despite unstable depth estimates of existing eye tracking systems. In addition, we propose an integrated learning process that leverages visual cues to help users learn how to voluntarily shift their focal depth. We also perform multiple user studies to evaluate user experience and mental model in our gaze-depth interaction process.
@@ 2.3. Immersive Gaze-based UI Design
There is a flurry of work on user interface design that incorporate gaze input in VR and AR systems. For example, Pfeuffer et al. (Pfeuffer et al., 2021; Blattgerste et al., 2018) proposes a design space for gaze-input based interfaces for AR applications. Hirzle et al. (Hirzle et al., 2019) introduced a design space for head mounted devices focusing on human depth perception and technical issues. In this paper, we propose a layer-based UI design for virtual reality that integrates active visual depth manipulation and gaze movement for VR interaction. We fully evaluate this new interaction and demonstrate the potential of this gaze-only interaction method as an effortless and intuitive method for everyday VR through a learning process.
@@ 3.2. Experiment Design For Pilot Study
[…]
@@ 4. FocusFlow: System Overview
  • Gaze-depth Interaction Design: FocusFlow leverages the proposed layer-based UI design and offers a gaze-only interaction mechanism by combining gaze direction and gaze depth. The gaze direction is used for selection of the target object for interaction and gaze depth is used for activating a certain layer in the UI. Section LABEL: discusses a variety of applications that can benefit from this hands-free gaze-only interaction design.
  • Adaptive Interaction Learning Process: FocusFlow takes advantages of a series of visual cues to help users master this brand-new interaction. The visual cues are designed such that they provide users with depth perception and learning opportunities to practice, receive feedback, and master intuitive adjustment of their gaze depth for interaction. Section LABEL: describes two different learning strategies based on in-stages and adaptive visual cues.
@@ 8.4.1. Activation Time
We define ’activation time’ as the duration from the moment a user’s gaze targets an object to when the virtual window appears. Figure 16(a) and 16(b) illustrate the activation times across all selection attempts by Group 2 during two tests in the testing stage. It’s important to note that dark squares marked as ’N/A’ represent instances where participants didn’t complete the selection task within the 5-second limit. Remarkably, participants in Group 2 successfully completed the selection tasks in the no-guidance test setting. This achievement underscores their high proficiency with the gaze-depth interaction method, particularly when contrasted with the last several attempts shown in Figure 12 for the in-stage learning strategy. Additionally, their activation times in the no-guidance test are comparable to the time in the adaptive-guidance test, suggesting that Group 2 participants have effectively mastered the gaze-depth interaction method without any visual guidance. Moreover, we also compare the activation time between the gaze-depth method (FocusFlow) and the baseline dwell method in Figure 18a. We could find that the two settings in FocusFlow achieve a similar activation time of around 1.3s, which is shorter than the 2-second dwell method, while the half-second dwell method achieved the shortest time.
@@ 9. Case Study
10. Discussion and Future Work
10.1. User Fatigue and Cognitive Load
Our user studies demonstrate the effectiveness of visual cues combined with the learning process to master this new interaction input and eventually use the system without any visual cues. However, some users still appreciate the scaffolding that these cues provide and showed interest in using the cues as a continual depth perception guideline. Therefore, visual cues are not merely transitional aids and can be an integral aspect of the everyday usage. This is analogous to television subtitles, where viewers might prefer the use of subtitles despite their capability in comprehending the show without them as a way of reducing cognitive load and bolster comprehension.
10.2. Compatibility and Extendibility

### v3

@@ Abstract.
Gaze interaction presents a promising avenue in Virtual Reality (VR) due to its intuitive and efficient user experience. Yet, the depth control inherent in our visual system remains underutilized in current methods. In this study, we introduce FocusFlow, a hands-free interaction method that capitalizes on human visual depth perception within the 3D scenes of Virtual Reality. We first develop a binocular visual depth detection algorithm to understand eye input characteristics. We then propose a layer-based user interface and introduce the concept of “Virtual Window” that offers an intuitive and robust gaze-depth VR interaction, despite the constraints of visual depth accuracy and precision spatially at further distances. Finally, to help novice users actively manipulate their visual depth, we propose two learning strategies that use different visual cues to help users master visual depth control. Our user studies on 24 participants demonstrate the usability of our proposed virtual window concept as a gaze-depth interaction method. In addition, our findings reveal that the user experience can be enhanced through an effective learning process with adaptive visual cues, helping users to develop muscle memory for this brand-new input mechanism. We conclude the paper by discussing potential future research topics of gaze-depth interaction.
Gaze Interaction; Visual Depth; Virtual Reality; 3D User Interface; Eye Tracking
journalyear: 2024copyright: acmlicensedconference: Proceedings of the CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, HI, USAbooktitle: Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613904.3642589isbn: 979-8-4007-0330-0/24/05ccs: Human-centered computing User interface designccs: Human-centered computing Interaction techniquesccs: Human-centered computing Graphical user interfacesccs: Human-centered computing Virtual reality
@@ 1. Introduction
Gaze interaction has gained popularity as an input method for 3D interaction in mixed reality (XR) headsets (Sidenmark et al., 2021; Lu et al., 2021a; Adhanom et al., 2023), where the eye movements and gestures would be used as an input. Current gaze-based XR interaction methods often rely on gaze as a means of pointing, supplemented by other inputs such as hand gestures (Yu et al., 2021), head movement (Sidenmark et al., 2021), or dwell time (Penkar et al., 2012; Piumsomboon et al., 2017) for selection confirmation to avoid the so-called ”Midas Touch” problem (Velichkovsky et al., 1997) – the problem of distinguishing intentional gaze inputs from involuntary fixations as the users perceive the scene. However, these works primarily utilize the direction of gaze and overlook the valuable visual depth information, representing an additional free input dimension along the z-axis (Figure 1). To see how, imagine a window through which you can see scenery outside. By bringing your gaze focus closer to the window, you blur out the distant scenery and are able to perceive dust particles on the window surface instead. Another example is autostereograms (Tyler and Clarke, 1990), where slightly adjusting your visual depth allows you to perceive a 3D image (see example in Appendix A). In both of these examples, users actively manipulate their visual depth to select which object to observe. In this research, we aim to leverage this natural eye behavior and convert it into an intuitive and learnable interaction input in Virtual Reality (VR).
@@
More recent works on gaze-based VR/AR interaction demonstrate the great potential of visual depth as an interaction input to solve the Midas touch problem. These methods either guide the user to look at physical or virtual objects at different depths (Ahn et al., 2020; Pai et al., 2016; Wang et al., 2022a, b; Sidenmark et al., 2023) or rely on voluntary eye convergence and divergence (Kirst and Bulling, 2016; Kudo et al., 2013) by asking the users focusing on the nose or imagining to fixate on some point behind the display plane. These works prove the feasibility of leveraging visual depth as an input for VR headset, however, the usability and learnability of this new interaction design are not fully explored. Failure-prone and difficult-to-learn experience can lead to user frustration and fatigue, making the interaction unusable. How to guide the users to learn to manipulate their visual depth as a reliable input is still a research gap.
@@
In particular, leveraging visual depth as an interaction input is challenging due to several practical factors. First, the natural eye movement is inherently unstable, resulting in noisy visual depth estimates. Moreover, the eye tracking modules in the headsets suffer from limited frame rate and jitters, adding extra noise to depth estimates. So, a usable interaction mechanism is required that is robust against imprecise eye input signal. Second, user’s visual depth fixation naturally relies on the presence of objects. So, it may feel counter-intuitive if the user is asked to transition to a depth without any visual referent. So, it is crucial to ensure that novice users can intuitively and effortlessly learn and use this interaction and a carefully designed learning procedure is needed to guide the users on how to master this new interaction mechanism. To tackle these challenges, we employed an empirical research approach. Initially, we devised a visual depth detection method for off-the-shelf VR headsets and conducted experiments to investigate the properties of visual depth input. Building upon the insights gained from these experiments, we subsequently developed our proposed user interface and incorporated a learning procedure with visual cues to help users adapt to the depth-control process.
As such, we present FocusFlow, a novel gaze-depth interaction method that realizes the concept of Virtual Window (just like the examples mentioned above) by utilizing the 3D nature of immersive scenes and the innate human capability to shift focus across different visual planes. FocusFlow employs the eye trackers in VR headsets to estimate the user’s visual depth. It then proposes a layer-based user interface, where each layer acts as a transparent Virtual Window positioned at varying depths. These windows become interactive when a user’s visual depth aligns with their location, allowing for seamless interaction without visual clutter. Specifically, FocusFlow selects interaction targets based on the user’s gaze direction and activates Virtual Windows—initially hidden within the VR scene—through voluntary changes in visual depth (as illustrated in Figure 1). This mechanism facilitates a range of applications, from quick previews and system panel activation to detailed zoom lenses for accessibility, enhancing the VR experience without overwhelming the user interface.
@@
To help users master this new interaction mechanism, we incorporate visual cues in the UI design and a learning process that allows users to establish a clear connection between the eye behavior and visual depth control, facilitating an intuitive and effortless interaction for novice users. The key intuition is that users can develop muscle memory for freely shifting their focus between different visual planes even if empty of objects. We propose two learning strategies, in-stages learning, and adaptive learning. The in-stages learning uses a series of strong to weak visual cues at a certain depth (e.g. where the transparent virtual window is located) so the user can practice visual depth control by just moving her gaze from the target object to the cue. While these cues are effective in practicing visual depth control, they also cause visual interference for the user during non-interaction times. The proposed adaptive visual cues addresses this issue by adaptively changing the transparency of visual cues as the user gets closer to the intended visual depth. The transparency level of the cue also serves as a feedback for the user to better control their visual depth changes. To help master this interaction, we design the adaptive learning process gradually reduces the guided depth range until reaching no visual cue.
@@
To evaluate the usability and learnability of FocusFlow, we conducted quantitative and qualitative evaluations, considering input efficiency, accuracy, and cognitive load. Results from 24 participants indicate that FocusFlow offers an efficient, easy-to-learn experience for hands-free selection tasks, effectively addressing the Midas Touch problem while maintaining natural interaction. Furthermore, our findings suggest that the adaptive visual cues in our design enhance users’ depth perception during visual depth transitions. Additionally, the adaptive learning strategy improves users’ proficiency with our gaze-depth interaction method.
The contribution of this work comprise:
@@
    We analyze the characteristics of gaze-depth estimates and propose strategies for utilizing this imprecise input information for interaction design.
@@
    We develop a layer-based user interface for gaze-depth interaction by realizing the concept of the “Virtual Window”. This interface supports intuitive and robust hands-free selection in VR while avoiding Midas Touch problem.
@@
    We design three visual cues that provide users with depth perception, and two learning strategies, in-stages learning and adaptive learning, to guide novice users in adapting to depth control and developing muscle memory for gaze-depth interaction.
@@
    We conduct a user study and a case study to assess both usability and learnability of our proposed system, and compare the effectiveness for different learning strategies.
@@ 2. Related Work
In this section, we first summarize different eye behaviors and the corresponding input techniques which can be used for gaze interactions. Based on this, we then review the evolution of gaze-based interaction in VR and some representative methods to identify the research gap. Lastly, we discussed design space research on user interfaces that integrate gaze input.
@@ 2.1. Eye Behaviors and Input Techniques
2.2. Gaze-Based VR Interaction
@@
Gaze-depth-based Interaction. Gaze vergence is an emerging eye-input information that has been used for VR interaction. Compared to multi-modal gaze interaction methods that leverage handheld controllers or hand gestures, visual depth offers a completely hands-free interaction mechanism. Previous works have shown the feasibility of visual depth estimations in the VR headsets by using binocular disparity and stereoscopic vergence in VR headset (Hirzle et al., 2019; Vidal et al., 2014). The concept of visual depth as a new interaction input has explored in the previous works, either by defining a semi-transparent window at a different focal depth (Pai et al., 2016; Wang et al., 2022a, b), tracking voluntary vergence movements (Kirst and Bulling, 2016; Ahn et al., 2020), or matching the vergence changes with the depth changes of a moving objects (Sidenmark et al., 2023) in VR. These studies have confirmed the efficacy of the gaze-depth-based interaction method from a technical standpoint. However, the learnability of this interaction poses a significant concern, as most users are not acquainted with gaze-depth input. There is a clear need for specific designs that instruct users on how to actively adjust their visual depth, along with comprehensive user studies to assess the usability and efficiency of this novel interaction mechanism. This paper targets these limitations and proposes a layer-based UI design to leverage active visual depth manipulation despite unstable depth estimates of existing eye tracking systems. In addition, we highlight an integrated learning process that leverages visual cues to help users learn how to voluntarily shift their focal depth. We also perform a user study and a case study to evaluate user experience and mental model in our gaze-depth interaction process.
@@ 2.3. Immersive Gaze-based UI Design
There is a flurry of work on user interface design that incorporate gaze input in VR and AR systems. For example, Pfeuffer et al. (Pfeuffer et al., 2021; Blattgerste et al., 2018) proposes a design space for gaze-input based interfaces for AR applications. Hirzle et al. (Hirzle et al., 2019) introduced a design space for head mounted devices focusing on human depth perception and technical issues. In this paper, we propose a layer-based UI design for virtual reality that integrates active visual depth manipulation and gaze movement for VR interaction. We evaluate this new interaction and demonstrate the potential of this gaze-only interaction method as an effortless and intuitive method for everyday VR through a learning process.
@@ 3.2. Experiment Design For Pilot Study
For the apparatus, we use an HTC VIVE Pro Eye headset, which has an FOV of 100 and a 90Hz frame rate. The eye tracking data is recorded by a built-in eye tracker camera at a frame rate of 120Hz and 0.5-1.1 tracking precision. To evaluate the accuracy of visual depth estimates, we invited 7 participants from the campus. For each test, the participant is sitting on a chair while wearing the headset and follows the instructions on looking at certain targets.
@@ 4. FocusFlow: System Overview
  • Gaze-depth Interaction Design: FocusFlow leverages the proposed layer-based UI design and offers a gaze-only interaction mechanism by combining gaze direction and gaze depth. The gaze direction is used for selection of the target object for interaction and gaze depth is used for activating a certain layer in the UI. Section 6.1 discusses a variety of applications that can benefit from this hands-free gaze-only interaction design.
  • Adaptive Interaction Learning Process: FocusFlow takes advantages of a series of visual cues to help users master this brand-new interaction. The visual cues are designed such that they provide users with depth perception and learning opportunities to practice, receive feedback, and master intuitive adjustment of their gaze depth for interaction. Section 7 describes two different learning strategies based on in-stages and adaptive visual cues.
@@ 8.4.1. Activation Time
We define ’activation time’ as the duration from the moment a user’s gaze targets an object to when the virtual window appears. Figure 17(a) and 17(b) illustrate the activation times across all selection attempts by Group 2 during two tests in the testing stage. It’s important to note that dark squares marked as ’N/A’ represent instances where participants didn’t complete the selection task within the 5-second limit. Remarkably, participants in Group 2 successfully completed the selection tasks in the no-guidance test setting. This achievement underscores their high proficiency with the gaze-depth interaction method, particularly when contrasted with the last several attempts shown in Figure 12 for the in-stage learning strategy. Additionally, their activation times in the no-guidance test are comparable to the time in the adaptive-guidance test, suggesting that Group 2 participants have effectively mastered the gaze-depth interaction method without any visual guidance. Moreover, we also compare the activation time between the gaze-depth method (FocusFlow) and the baseline dwell method in Figure 18a. We could find that the two settings in FocusFlow achieve a similar activation time of around 1.3s, which is shorter than the 2-second dwell method, while the half-second dwell method achieved the shortest time.
@@ 9. Case Study
10. Limitation and Future Work
FocusFlow features a hands-free gaze-depth-based input method that is efficient and intuitive for selection tasks after a minimal learning process. We proposed a layer-based user interface design according to the visual depth estimation characteristics and learning strategies to help users learn this gaze-depth input method. Although our evaluation results demonstrate FocusFlow is superior to baseline in terms of operational efficiency and the effectiveness of the adaptive learning strategy in helping users master this gaze-depth input method, there are still some aspects that could be improved or extended in future works:
Statistical Analysis: The current paper does not include any major statistical significance analysis due to small participant population. In our future work, we plan to perform a long-term experiment with a larger user population to fully demonstrate the effectiveness and intuitiveness of the proposed approach.